OpenStudy (anonymous):

What number is 95% of 350?

OpenStudy (anonymous):

Have you set up a proportion? x/350 = 95/100

OpenStudy (campbell_st):

well 10% = 35 find 5% by halving the value of 10% the subtract it from 350. hope it helps

OpenStudy (anonymous):

So then cross multiply. 100x = 95*350

OpenStudy (anonymous):

then divide both sides for 100 to solve for x.

OpenStudy (anonymous):

wait so it is 100 over x or x over 100, and is it 95 over 350 or 350 over 95... sorry I suck at math

OpenStudy (anonymous):

Make sense? I find that the easiest and most straightforward to do it. I just think what number is 95 percent of 350 translates nicely to what number divided by 350 is equal to the ratio 95/100(95 percent)

OpenStudy (anonymous):

x over 350 = 95/100

OpenStudy (anonymous):

x/350=95/100 aka x over 350 = 95 over 100

OpenStudy (campbell_st):

this is the simple non calculator method 10% move the decimal point 1 place left so 10% = 35 now 5% = 10%/2 or 35/2 = 17.5 subtract it from 350 (100%) to get 95%

OpenStudy (anonymous):

Then from there you can solve algebreically.

OpenStudy (anonymous):

So Campbell_st is the answer 17.5 your saying?

OpenStudy (anonymous):

Cora is graphing the equation . She has placed a point on the origin, and uses the slope to find 2 more points. Which points does Cora plot on her graph? Choose exactly two answers that are correct. A. (−1, −12) B. (−12, −1) C. (1, 12) D. (12, 1)

OpenStudy (anonymous):

c and a?

OpenStudy (campbell_st):

no the answer is 100% - 5 % = 95% so 350 - 17.5 = ?

OpenStudy (campbell_st):

? being the answer

OpenStudy (anonymous):

THX

OpenStudy (anonymous):

I got 332.5